OpenStudy (anonymous):

Help please, Find the vertex,focus and directrix of the following parabola. (x+3)^2=-3(y-2)

OpenStudy (bahrom7893):

(x+3)^2=-3(y-2) (x-(-3))^2 = -3y + 6 3y = -1*(x-(-3))^2 + 6

OpenStudy (bahrom7893):

I'm just trying to get it into the regular form: y = (-1/3)*(x-(-3))^2 + 2

OpenStudy (anonymous):

okay...i got (-3,2) for vertex for focus i got (-3,-3) and for directrix i got -1 are these correct?

OpenStudy (bahrom7893):

y-2 = (-1/3)*(x-(-3))^2 vertex is (h,k), so in this case: (-3, 2). yes vertex is right

OpenStudy (bahrom7893):

a = -1/3, h = -3, k = 2

OpenStudy (bahrom7893):

focus is (a + h, k) ((-1/3)-3, 2) = (-10/3, 2)

OpenStudy (anonymous):

okay i see

OpenStudy (bahrom7893):

what was the formula for directrix? amistre am i doing this right?

OpenStudy (bahrom7893):

I dont remember any of this, im looking up formulas that might be wrong.

OpenStudy (amistre64):

directix is the opposite direction, same distance, as focus

OpenStudy (amistre64):

y^2 = 4ax is a standard parabola ; where "a" measure is the distance from vertex to focus/directix

OpenStudy (amistre64):

(x+3)^2=-3(y-2) x^2 = -3y is our parabola centerd at the origin so that we can see it better -3 = 4a a = -3/4

OpenStudy (anonymous):

so that is the directrix ?..also does the mean the focus was wrong?

OpenStudy (anonymous):

*this

OpenStudy (amistre64):

it means that the distance to our focus is -3/4 from the vertex; and the directix is in the other direction

OpenStudy (anonymous):

ahhh okay i understand the picture helped

OpenStudy (amistre64):

our vertex is at -3,2 right?

OpenStudy (amistre64):

-3,-6 might be better

OpenStudy (amistre64):

(x+3)^2 = -3y + 6 (x+3)^2 - 6= -3y -1/3 (x+3)^2 - 6= y yeah, vertex = -3,-6 our focus is then at (-3,(-6- 3/4)); and the directix is the horisontal line at y=(-6+3/4)

OpenStudy (amistre64):

almost did that right lol

OpenStudy (anonymous):

yes..okay i understand you..makes sense now thanks..amistre and bahrom! :D

OpenStudy (amistre64):

if you can find my mistake, and correct it, itll prolly be helpful

OpenStudy (amistre64):

at any rate, click on properties and itll help you chk your answers http://www.wolframalpha.com/input/?i=%28x%2B3%29%5E2%3D-3%28y-2%29

OpenStudy (anonymous):

Okay will do thanks